# Can a Quantum Computer Simulate Nuclear Magnetic Resonance Spectra Better than a Classical One?

### v1

@@ Abstract
The simulation of the spectra measured in nuclear magnetic resonance (NMR) spectroscopy experiments is a computationally non-trivial problem, and as such, it represents a problem for which a quantum computer may provide some practical advantage over traditional computing methods. In order to understand the extent to which such problems may provide examples of useful quantum advantage, it is important to understand the limitations of existing classical simulation methods. In this work, we benchmark our classical solver designed to solve such problems. We find that it performs well, even beyond the common experimental parameter regimes, except for a specific molecule with certain unusual features. We discuss what implications this may have for future efforts to demonstrate quantum advantage in the context of NMR.
@@ I Introduction
While advances continue to be made in the field of quantum computing hardware [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11], demonstrating concrete examples of useful quantum advantage remains a challenge [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33]. Considerable effort has been invested in searching for problems in which a solution provided by a quantum computer presents some advantage over a solution provided by a classical computer (including the extreme case in which a classical computer cannot reasonably provide a solution at all). One such problem which has been proposed as a candidate for demonstrating quantum advantage is the prediction of the spectra measured in nuclear magnetic resonance (NMR) spectroscopy experiments. Such a problem is seen as a natural candidate for quantum computation, as the Hamiltonian describing a molecule studied through NMR takes the form of a typical quantum spin system, the degrees of freedom of which map in a straightforward fashion to the qubits of a quantum computer. Previous work [34, 35, 36, 37] has investigated the manner in which such a problem can be solved through the use of a quantum computer.
@@
Of course, understanding whether such a problem is truly capable of demonstrating useful quantum advantage requires understanding the limits of classical computation in the context of computing NMR spectra. In [38, 39, 40, 41, 42, 43, 44] the extent to which various classical approximation methods can tackle such a problem has been explored. In this work, we propose our own classical method for simulating NMR spectra, and examine its performance when applied to a selection of real-world molecules. We find that this solver provides accurate results in essentially all experimentally realistic parameter regimes, except for a small class of molecules with particularly unusual properties. We use the performance of our solver applied to these molecules to gain insight into the types of problems for which we may hope to see some degree of quantum advantage in the context of NMR spectroscopy.
@@
The outline of this paper is as follows: In Section II we briefly discuss the basics of NMR spectroscopy, and in Section III we explain the details of our solver. In Section IV we discuss the performance of our solver when applied to a collection of molecules for which we possess an exact solution, and in Section IV we consider this performance for molecules for which we do not have an exact solution. We conclude in Section V. In Appendix A we explain how the particular spectral function we compute is related to quantities measured in an NMR experiment. In Appendix B we discuss how we compute certain accuracy metrics based on the output of our solver. In Appendix C we list where we obtain the parameters which define the molecules we study in this work. In Appendix D we display some additional supporting data which was omitted from the main text.
@@ II Nuclear Magnetic Resonance
Nuclear magnetic resonance spectroscopy is an important tool in analytical chemistry, used to identify molecules and to elucidate their structure, among other applications [45, 46, 47, 48, 49, 50, 51]. In broad terms, NMR spectroscopy involves placing a sample of a given molecule in a large, constant magnetic field, perturbing it with a electromagnetic pulse, and measuring the response. In this section we provide a very brief discussion of the quantity which is measured in a typical NMR experiment, but in enough detail to elucidate the basic calculations which must be performed by our solver. We assume that the molecules are in a liquid solution, which simplifies some aspects of the analysis due to motional averaging.
@@
$\hat{H}=-\sum_{l}\gamma_{l}\left(1+\delta_{l}\right)B^{z}\hat{I}^{z}_{l}+2\pi\sum_{k<l}J_{kl}\mathbf{\hat{I}}_{k}\cdot\mathbf{\hat{I}}_{l}$ (1)
The objects $\hat{I}^{\alpha}=\hat{S}^{\alpha}/\hbar$ with $\alpha\in\{x,y,z\}$ correspond to the usual SU(2) spin operators for a spin-S particle, which represent atomic nuclei present in the molecule. Only atomic nuclei with non-zero nuclear spin (“active nuclei”) will be relevant for an NMR experiment, with focus often being placed on nuclei with spin-1/2, such as 1H or 31P nuclei. The molecule is subjected to a magnetic field which we take to be in the Z direction, denoted by $B^{z}$, and each nuclei interacts with this field through a Zeeman interaction described by the gyromagnetic ratio $\gamma$. The different atomic nuclei can also couple to each other, parameterized by the terms $J_{kl}$. While each $J_{kl}$ could in principle be a tensor quantity, the motional averaging occurring in solution results in a single scalar quantity, thus resulting in an isotropic Heisenberg coupling. This term is generally a result of indirect spin-spin coupling mediated by electrons in chemical bonds.
@@
allows for the identification of different chemical groups present in the molecule. In proton NMR, which examines the NMR spectra of molecules that result primarily from the interaction of 1H nuclei with the external magnetic field, we define $\delta=0$ according to the effective gyromagnetic ratios experienced by the Hydrogen nuclei in the molecule tetramethylsilane (TMS), therefore setting the value of $\gamma$, as well as the values of the chemical shifts for the nuclei in any other given molecule (these gyromagnetic ratios in TMS happen to all be identical).
When a sample is placed in an NMR spectrometer, the magnetic field of the device results in some net magnetization of the sample. A brief electromagnetic pulse flips the spins into the XY plane, which therefore begin to precess, creating a signal that can be recorded by measuring the induced magnetic field along the X or Y axis. In Appendix A, we show that the measured X and Y components of this magnetic field directly lead to the so-called spectral function, which, up to proportionality, is defined as,
$C\left(\omega\right)\propto\eta\sum_{n,m}\frac{\langle E_{n}|M^{-}|E_{m}\rangle\langle E_{m}|M^{+}|E_{n}\rangle}{\eta^{2}+[\omega-(E_{n}-E_{m})]^{2}}$ (3)
@@
$M^{\pm}\equiv\sum_{i}\gamma_{i}\left(\hat{I}^{x}_{i}\pm i\hat{I}^{y}_{i}\right)\equiv\sum_{i}\gamma_{i}\hat{I}^{\pm}_{i}\,.$ (4)
For the applicability of Eq. (3) see Appendix A. Note that we are typically not interested in the overall normalization of the spectral function, so in this work we will always normalize this function so that its integral is equal to the number of active nuclei in the molecule.
This (real and positive) quantity consists of a collection of Lorentzian peaks, centered at the transition frequencies of the system. The amplitudes of these peaks are determined by the matrix elements of the $M^{\pm}$ operators, which are restricted by the fact that the NMR Hamiltonian conserves the total spin along the Z axis,
@@
This conservation allows the Hamiltonian to be written in a block-diagonal form, such that the largest block has dimension $\mathcal{D}^{2}$, with
$\mathcal{D}={\textstyle\binom{N}{N/2}}\;\sim\mathcal{O}\left((2^{N}/\sqrt{N})\right)\,$, (6)
where $N$ is the number of active nuclei in the molecule. If we were to attempt to calculate the spectral function directly by performing an exact diagonalization of the Hamiltonian on a classical computer, we would therefore anticipate that the required memory would scale as $\mathcal{O}\left(2^{2N}/N\right)$, while the required computational time would scale as $\mathcal{O}\left(2^{3N}/N^{3/2}\right)$. For specific molecules with highly symmetric coupling, one can additionally exploit a local SU(2) symmetry to reduce the effective system size [52]. For example, a freely rotating methyl group can be replaced by a decomposition into a spin-3/2 particle and two spin-1/2 particles, effectively decreasing the system size by one spin. However, the combinatorial effort associated with this reduction increases the complexity of the problem, and for small molecules (with less than around twelve nuclei), this can lead to an overall run time which is in fact longer. Furthermore, for molecules beyond a size of approximately twenty spins, neither of these symmetries is sufficient to obtain an exact solution in a reasonable amount of computational time, and further approximations must therefore be made. These approximations will be the basis for our solver, the details of which we discuss in the next section.
@@
The quantity $\eta$, which we refer to as the “broadening” (sometimes known as the “line width”), is a convergence factor which is meant to model the rate at which the measured signal decays due to decoherence processes and to model a finite spectrometer resolution. In the non-interacting case, it is a straightforward exercise to verify that there is one peak corresponding to each nucleus, centered at the Larmor frequency of that nucleus. This quantity is therefore extremely useful in determining the chemical shifts of the molecule in question. In the presence of interactions, these peaks will typically split apart into additional smaller peaks, as the interaction terms delocalize the eigenstates of the system and allow for additional transitions, at slightly shifted frequencies. Therefore, it is generally desirable to operate NMR spectrometers at the largest possible magnetic field, since higher magnetic fields will enhance the effect of the Zeeman interaction compared with both the Heisenberg couplings and the broadening scale, leading to an NMR spectrum which can more easily identify the chemical shifts of the constituent nuclei. In this work, we will mostly focus on proton NMR. The parameter regimes in which typical NMR experiments operate result in a reference Larmor frequency $\omega_{\text{ref}}$ for the 1H nuclei in TMS which corresponds to a natural frequency $\nu_{\text{ref}}=\omega_{\text{ref}}/2\pi$ of a few hundred MHz, although smaller benchtop devices typically operate at lower frequencies, potentially reaching as high as 80 MHz. The chemical shifts in a typical molecule examined in proton NMR usually range from around $10^{-6}$ to $10^{-5}$ (often quoted as 1 to 10 “parts per million,” or “ppm” for short), resulting in peaks in the spectral function which are separated by anywhere from a few hundred to a few thousand Hz, while the coupling terms are typically on the order of tens of Hz. A broadening of a few Hz is usually reasonable. In our analysis in this paper, we will study three magnetic field strengths, corresponding to Larmor frequencies of 400, 80, and 20 MHz, which, for the purposes of this work, we will refer to as high, low, and very low field. We will also consider two different broadening values, corresponding to Lorentzian peaks with a full-width at half maximum of 1 Hz, and also 0.1 Hz, which we refer to as high and low broadening. Note that the parameter regimes corresponding to very low field or low broadening are somewhat atypical with respect to actual experiments and merely serve as a stress test for our solver.
@@ III The Spin-Dependent Cluster Solver
To make a prediction for the spectral function which will be measured in an NMR experiment, we must compute the spectral function in equation (3). A brute-force solution to this problem would require either performing a time evolution of the Hamiltonian dynamics and then a transformation into the frequency domain, or an exact diagonalization of the Hamiltonian, which can be used to compute the sum in equation (3) directly. Due to the exponential scaling of the Hilbert space, such a brute-force approach will become very computationally expensive after about twenty spins, and computationally prohibitive after a few more, even if we do make use of the conservation of total spin along the Z axis and the SU(2) symmetry of symmetrically coupled groups.
To circumvent this issue, our solver is designed to compute the “spin-resolved” quantity
$C_{i}\left(\omega\right)=\gamma_{i}\sum_{j}\gamma_{j}\eta\sum_{n,m}\frac{\langle E^{(\Gamma_{i})}_{n}|\hat{I}^{-}_{i}|E^{(\Gamma_{i})}_{m}\rangle\langle E^{(\Gamma_{i})}_{m}|\hat{I}^{+}_{j}|E^{(\Gamma_{i})}_{n}\rangle}{\eta^{2}+[\omega-(E^{(\Gamma_{i})}_{n}-E^{(\Gamma_{i})}_{m})]^{2}}$ (7)
@@
$C\left(\omega\right)\approx\sum_{i}C_{i}\left(\omega\right).$ (8)
Dividing up the computation in this way, where we compute each $C_{i}$ separately, is advantageous because it allows us to take advantage of the key approximation utilized in the NMR solver, which is to construct a cluster $\Gamma_{i}$ of spins which are most likely to influence the physics of the chosen $C_{i}$, and then diagonalize the reduced Hamiltonian of this cluster, disregarding the effects from spins outside of the cluster. This approximation becomes exact in the limit where the cluster size coincides with the number of active nuclei. Note that these clusters can overlap, which is essential for obtaining accurate results. Furthermore, at fixed cluster size, this approach leads to a linear scaling with total molecular system size, though additional runtime improvements can further be made by diagonalizing the Hamiltonian for spins located in identical clusters only once, and also by exploiting the symmetry considerations outlined above, for each cluster Hamiltonian independently.

### v2

@@ Abstract
The simulation of the spectra measured in nuclear magnetic resonance (NMR) spectroscopy experiments is a computationally non-trivial problem which, due to its natural interpretation as a quantum spin problem, maps in a straightforward way to a quantum computer. As such, it represents a problem for which such a device may provide some practical advantage over traditional computing methods. In order to understand the extent to which such problems may indeed provide examples of useful quantum advantage, it is important to understand the limitations of existing classical simulation methods. In this work, we benchmark our own classical solver designed to study such problems. This solver uses a clustering approximation to achieve a resource scaling which is linear in the total number of nuclear spins in a given molecule, for a fixed cluster size. The success of such an approximation would present a stark repudiation to the common claim that such problems require an exponential scaling of resources, the very claim which makes simulating an NMR spectra a candidate for quantum advantage. Our benchmarking results indicate that our approximation performs well throughout, and even somewhat beyond, the more typical experimental regimes. We discuss what implications this may have for future efforts to demonstrate quantum advantage in the context of NMR.
@@ I Introduction
While advances continue to be made in the field of quantum computing hardware [31, 32, 57, 47, 18, 24, 38, 8, 19, 33, 50], demonstrating concrete examples of useful quantum advantage remains a challenge [46, 54, 55, 17, 49, 27, 45, 62, 10, 41, 11, 16, 1, 42, 6, 5, 20, 3, 44, 23, 4, 30]. Considerable effort has been invested in searching for problems in which a solution provided by a quantum computer presents some advantage over a solution provided by a classical computer (including the extreme case in which a classical computer cannot reasonably provide a solution at all). One such problem which has been proposed as a candidate for demonstrating quantum advantage is the prediction of the spectra measured in nuclear magnetic resonance (NMR) spectroscopy experiments. Such a problem is seen as a natural candidate for quantum computation, as the Hamiltonian describing a molecule studied through NMR takes the form of a typical quantum spin system, the degrees of freedom of which map in a straightforward fashion to the qubits of a quantum computer. Previous work [53, 52, 29, 48, 9] has investigated the manner in which such a problem can be solved through the use of a quantum computer.
@@
Of course, understanding whether such a problem is truly capable of demonstrating useful quantum advantage requires understanding the limits of classical computation in the context of computing NMR spectra. In [34, 35, 14, 36, 37, 25, 56] the extent to which various classical approximation methods can tackle such a problem has been explored. In this work, we propose our own classical method for simulating NMR spectra, and examine its performance when applied to a selection of real-world molecules. This solver makes use of a cluster approximation which is designed so that for a fixed cluster size its resource consumption is inherently linear in the total number of nuclear spins in the molecule. We find that this approximation provides accurate results in the experimental regimes which are typical for standard NMR experiments. Our benchmarking analysis does not reveal any molecules which require a cluster size of more than a dozen or so spins in order to achieve reasonably accurate spectra. These results suggest that the problem of computing accurate NMR spectra in the standard experimental regimes is tractable with resources that scale linearly with the number of nuclear spins. We also consider regimes with magnetic fields and linewidths that are somewhat lower than what is typical in standard NMR spectrometers, and find good performance here as well. We use the convergence of our cluster approximation applied to these molecules to gain insight into the types of problems for which we may hope to see some degree of quantum advantage in the context of NMR spectroscopy.
@@
The outline of this paper is as follows: In Section II we briefly discuss the basics of NMR spectroscopy, and in Section III we explain the details of our solver. In Section IV we discuss the performance of our solver when applied to a collection of molecules for which we possess an exact solution, and in Section V we consider this performance for molecules for which we do not have an exact solution. In Section VI we discuss the necessary extensions to our basic solver when handling certain special molecules. We conclude in Section VII. In Appendix A we explain how the particular spectral function we compute is related to quantities measured in an NMR experiment. In Appendix B we discuss how we compute certain accuracy metrics based on the output of our solver. In Appendix C we list where we obtain the parameters which define the molecules we study in this work. In Appendix D we display some additional supporting data which was omitted from the main text.
@@ II Nuclear Magnetic Resonance
Nuclear magnetic resonance spectroscopy is an important tool in analytical chemistry, used to identify molecules and to elucidate their structure, among other applications [7, 51, 15, 60, 13, 39, 22, 28, 12]. In broad terms, NMR spectroscopy involves placing a sample of a given molecule in a large, constant magnetic field, perturbing it with an electromagnetic pulse, and measuring the response. In this section we provide a very brief discussion of the quantity which is measured in a typical NMR experiment, but in enough detail to elucidate the basic calculations which must be performed by our solver. We assume that the molecules are in a liquid solution, which simplifies some aspects of the analysis due to motional averaging.
@@
$\hat{\cal{H}}=-\sum_{l}\gamma_{l}\left(1+\delta_{l}\right)B^{z}\hat{I}^{z}_{l}+2\pi\sum_{k<l}J_{kl}\mathbf{\hat{I}}_{k}\cdot\mathbf{\hat{I}}_{l}$ (1)
The objects $\hat{I}^{\alpha}=\hat{S}^{\alpha}/\hbar$ with $\alpha\in\{x,y,z\}$ correspond to the usual SU(2) spin operators for a spin-S particle, which represent atomic nuclei present in the molecule. Only atomic nuclei with non-zero nuclear spin (“active nuclei”) will be relevant for an NMR experiment, with focus often being placed on nuclei with spin-1/2, such as 1H or 31P nuclei. The molecule is subjected to a magnetic field which we take to be in the Z direction, denoted by $B^{z}$, and each nucleus interacts with this field through a Zeeman interaction described by the gyromagnetic ratio $\gamma$. The different atomic nuclei can also couple to each other, parameterized by the terms $J_{kl}$. While each $J_{kl}$ could in principle be a tensor quantity, the motional averaging occurring in solution results in a single scalar quantity, thus resulting in an isotropic Heisenberg coupling. This term is generally a result of indirect spin-spin coupling mediated by electrons in chemical bonds.
@@
allows for the identification of different chemical groups present in the molecule. For each species of active nuclei, a standard reference compound is used to define $\delta=0$. For the case of 1H nuclei, for example, we make use of the convention that the chemical shifts of the hydrogen nuclei in the molecule tetramethylsilane (TMS) are defined as zero (these hydrogen nuclei in TMS all possess identical chemical shifts). This therefore sets the value of $\gamma$ for hydrogen nuclei in general, and allows for the assignment of chemical shift values for the hydrogen nuclei in any other given molecule.
When a sample is placed in an NMR spectrometer, the magnetic field of the device results in some net magnetization of the constituent nuclei. In a standard 1D NMR experiment, a brief electromagnetic pulse flips one species of active nuclei into the XY plane, which therefore begin to precess, creating a signal that can be recorded by measuring the induced magnetic field along the X or Y axis. In this work, we will focus exclusively on proton NMR, in which 1H nuclei are the active nuclei which are flipped into the XY plane. In Appendix A, we show that the measured X and Y components of this magnetic field directly lead to the so-called spectral function, which, up to proportionality, is defined as,
$C\left(\omega\right)\propto\eta\sum_{n,m}\frac{\langle E_{n}|\hat{M}^{-}|E_{m}\rangle\langle E_{m}|\hat{M}^{+}|E_{n}\rangle}{\eta^{2}+[\omega-(E_{n}-E_{m})]^{2}}$ (3)
@@
$\hat{M}^{\pm}\equiv\sum_{i}\gamma_{i}\left(\hat{I}^{x}_{i}\pm i\hat{I}^{y}_{i}\right)\equiv\sum_{i}\gamma_{i}\hat{I}^{\pm}_{i}\,.$ (4)
For the applicability of Eq. (3) see Appendix A. Note that we are typically not interested in the overall normalization of the spectral function, and since we are considering proton NMR, we will always normalize this function so that its integral is equal to the number of hydrogen nuclei in the molecule.
This (real and positive) quantity consists of a collection of Lorentzian peaks, centered at the transition frequencies of the system. The width of these peaks is controlled by the “broadening” $\eta$ (sometimes also known as the “line width”), which is formally a convergence factor, but can also be used to model a limited spectrometer resolution (which results from several experimental factors, including the decay of the measured signal due to decoherence). The amplitudes of these peaks are determined by the matrix elements of the $\hat{M}^{\pm}$ operators, which are restricted by the fact that the NMR Hamiltonian conserves the total spin along the Z axis,
@@
In the non-interacting case, it is a straightforward exercise to verify that there is one Lorentzian peak corresponding to each flipped nucleus, centered at the Larmor frequency of that nucleus. This quantity is therefore extremely useful in determining the chemical shifts of the molecule in question. In the presence of interactions, these peaks will typically split apart into additional smaller peaks, as the interaction terms delocalize the eigenstates of the system and allow for additional transitions, at slightly shifted frequencies. Therefore, it is generally desirable to operate NMR spectrometers at the largest possible magnetic field, since higher magnetic fields will enhance the effect of the Zeeman interaction compared with both the Heisenberg couplings and the broadening scale, leading to an NMR spectrum which can more easily identify the chemical shifts of the constituent nuclei.
The parameter regimes in which typical proton NMR experiments operate result in a natural frequency $\nu_{\text{ref}}=\omega_{\text{ref}}/2\pi$ of a few hundred MHz for the 1H nuclei in TMS, although smaller benchtop devices typically operate at lower frequencies, e.g. 80 MHz. The chemical shifts in a typical molecule examined in proton NMR usually range from around $10^{-6}$ to $10^{-5}$ (often quoted as 1 to 10 “parts per million,” or “ppm” for short), resulting in peaks in the spectral function which are separated by anywhere from a few hundred to a few thousand Hz, while the coupling terms are typically on the order of tens of Hz. A broadening of a few Hz is usually reasonable. In our analysis in this paper, we will study three magnetic field strengths, corresponding to natural frequencies of 400, 80, and 20 MHz, which, for the purposes of this work, we will refer to as high, low, and very low field. We will also consider two different broadening values, corresponding to Lorentzian peaks with a full-width at half maximum of 1 Hz, and also 0.1 Hz, which we refer to as high and low broadening. Note that the parameter regimes corresponding to very low field or low broadening are somewhat atypical with respect to actual experiments and merely serve as a stress test for our solver.
@@ III The Spin-Dependent Cluster Solver
To make a prediction for the spectral function which will be measured in an NMR experiment, we must compute the quantity in equation (3). A brute-force solution to this problem would require either performing a time evolution of the Hamiltonian dynamics and then a transformation into the frequency domain, or an exact diagonalization of the Hamiltonian, which can be used to compute the sum in equation (3) directly. The conservation of total spin along the Z axis allows the Hamiltonian to be written in a block-diagonal form, such that the largest block has size $\mathcal{D}\times\mathcal{D}$, with a reduced Hilbert space dimension of
$\mathcal{D}\sim{\textstyle\binom{N}{N/2}}\;\sim\mathcal{O}\left((2^{N}/\sqrt{N})\right)\,$, (6)
where $N$ is the number of active nuclei in the molecule. If we were to attempt an exact diagonalization of the Hamiltonian on a classical computer, we would therefore anticipate that the required memory would scale as $\mathcal{O}\left(2^{2N}/N\right)$, while the required computational time would scale as $\mathcal{O}\left(2^{3N}/N^{3/2}\right)$. For specific molecules with highly symmetric coupling, one can additionally exploit a local SU(2) symmetry to reduce the effective system size [58]. For example, a freely rotating methyl group can be replaced by a decomposition into a spin-3/2 particle and two spin-1/2 particles, effectively decreasing the system size by one spin. However, the combinatorial effort associated with this reduction increases the complexity of the problem, and for small molecules (with less than around twelve nuclei), this can lead to an overall run time which is in fact longer. Furthermore, for molecules beyond a size of approximately twenty spins, neither of these symmetries is sufficient to obtain an exact solution in a reasonable amount of computational time.
To circumvent this issue, our solver is designed to compute the “spin-resolved” spectral function
$C_{i}\left(\omega\right)=\gamma_{i}\sum_{j}\gamma_{j}\eta\sum_{n,m}\frac{\langle E_{n}|\hat{I}^{-}_{i}|E_{m}\rangle\langle E_{m}|\hat{I}^{+}_{j}|E_{n}\rangle}{\eta^{2}+[\omega-(E_{n}-E_{m})]^{2}}$ (7)
@@
$C\left(\omega\right)=\sum_{i}C_{i}\left(\omega\right).$ (8)
Dividing up the computation in this way, where each $C_{i}$ is computed separately, is advantageous because it allows us to take advantage of the key approximation utilized in the NMR solver, which is to construct a cluster $\Gamma_{i}$ of spins which are most likely to influence the physics of the chosen $C_{i}$. The solver diagonalizes the reduced Hamiltonian of this cluster, disregarding the effects from spins outside of the cluster, resulting in an approximation to the spin-resolved spectral function,
$C_{i}\left(\omega\right)\approx\gamma_{i}\sum_{j}\gamma_{j}\eta\sum_{n,m}\frac{\langle E^{(\Gamma_{i})}_{n}|\hat{I}^{-}_{i}|E^{(\Gamma_{i})}_{m}\rangle\langle E^{(\Gamma_{i})}_{m}|\hat{I}^{+}_{j}|E^{(\Gamma_{i})}_{n}\rangle}{\eta^{2}+[\omega-(E^{(\Gamma_{i})}_{n}-E^{(\Gamma_{i})}_{m})]^{2}}$ (9)
which uses the approximate energy eigenstates $|E^{(\Gamma_{i})}_{m}\rangle$ corresponding to the given cluster. This approximation becomes exact in the limit where the cluster size coincides with the number of active nuclei. Note that the clusters corresponding to different spins can overlap, which is essential for obtaining accurate results. Furthermore, at fixed cluster size, this approach leads to a linear scaling with total molecular system size, though additional runtime improvements can be made by diagonalizing the Hamiltonian for spins located in identical clusters only once, and also by exploiting the symmetry considerations discussed previously, for each cluster Hamiltonian independently.
@@
We emphasize that such a metric will have value zero for all spins that are not directly coupled, which may render it somewhat less useful for ordering the importance of spins in molecules which possess fewer direct couplings. However, as we will soon see, even this relatively simple metric results in a clustering approximation which is highly effective at computing the spectra of almost all of the molecules we study, and struggles only in a few special cases. For these special molecules, straightforward considerations regarding their structure lead us to a generalized clustering scheme which allows us to once again obtain accurate predictions, which we will discuss in more detail later.